# Atomic forces from correlation energy functionals based on the adiabatic- connection fluctuation-dissipation theorem

### v1

@@ IV Implementation and testing
Given the good quality found for our forces on this first test, we have then proceeded to determine their quantitative accuracy by comparing our analytical force calculations to the total energy finite-difference estimations. For this analysis, two different systems have been considered: the H2O molecule and the wurtzite phase of the boron nitride solid ($\omega$-BN, space group $P6_{3}/mmc$). For the H2O molecule, calculations were performed using a 120 Ry cutoff for the kinetic energy, and 20 bohrs for the simulation cell size. With this setting, the PBE structure and PBE vibrational frequencies were well converged when using the forces. For the RPA calculations, we used the same plane-wave cutoff and cell size. For the evaluation of the correlation components, we used 10 eigenvalues per valence electron ($N_{\rm eig}$=80), and 12 imaginary frequency points. Starting from the RPA equilibrium geometry (see Tab. I), we then varied one of the O-H bond length around 0.96 Å, in steps of 0.01 Å. The total energy finite-difference estimations were performed using the highly accurate five-point stencil derivation formula. For the $\omega$-BN solid, we first optimized the unit cell using the PBE functional and then displaced one B3+ cation along the $\mathbf{c}$ crystal axis. For these test calculations, we used a cutoff of 100 Ry, and a $\times2$$\times$2 k-point grid. For RPA, we used a grid made of 10 imaginary frequency points, determined 8 eigenvalues per valence electron ($N_{\rm eig}$=128), and used a $\times2$$\times$2 q-point grid. The error analyses for H2O and $\omega$-BN are presented in panel a) and b) of Fig. 3, respectively.
@@
In panel a), we note that the largest error on the force is found for PBE, with a value of $2\!\times\!10^{-4}$ Ry/bohr, substantially above any of the other approximations considered. This error reduces, however, by increasing the energy cutoff, either on the plane-waves or the density. At the level of HF, we instead note an almost perfect agreement between the analytical evaluation of the force and the numerical estimate from the total energy, with a difference between the two predictions below $10^{-6}$ Ry/bohr. The EXX method displays, at most, an error of $1.6\!\times\!10^{-5}$ Ry/bohr across the whole range of distances monitored. As discussed in Ref. [41], the oscillations and the overall error can be dampened further by increasing the kinetic energy cutoff and by tightening the self-consistent convergence conditions. If we compare the effect of self-consistency in the calculation, we see that the error found at the HF[PBE] level is larger than for HF and EXX. Despite this, the overall agreement remains of good quality, with a maximum error below $6.5\!\times\!10^{-5}$ Ry/bohr. A similar evolution is observed for the RPA method. At self-consistency, the error on RPA forces is found below $1\!\times\!10^{-5}$ Ry/bohr, similar to EXX. Without self-consistency, the evolution of RPA[PBE] matches HF[PBE], with a maximum error of $5.3\!\times\!10^{-5}$ Ry/bohr. Overall, the errors found are sufficiently small for exploiting the forces in a geometry relaxation procedure, given the 0.001 Ry/a0 default threshold set on each individual force within QE.
@@
Finally, in both panels a) and b), we note that if we omit the extra xc force terms for EXX and RPA, the methods denoted EXX $\{-{\bf F}^{\rm scf\text{-}EXX}_{\rm x}\}$ and RPA $\{-{\bf F}^{\rm scf\text{-}RPA}_{\rm x}\}$, respectively, the error in the analytical evaluation of the force is as large as 0.1-0.01 Ry/a0. This further confirms the importance of including these terms for any approximation that relies on the OEP method.
@@ V Results and discussion
Below we present our results for a set of molecules (H2, LiH, FH, N2, H2O, NH3, and CH4), and periodic solids (C, Si, and Ge).
@@ V.1 Molecules
We have evaluated the performance of the ACFDT-based methods with respect to HF, PBE, and reference CCSD(T) calculations on a set of covalently bonded molecules: H2, LiH, FH, N2, H2O, NH3, and CH4. For every system considered, a kinetic energy cutoff of 100 Ry has been used, and a cubic simulation cell of dimensions 20 bohrs has been used except for LiH, N2, and CH4, for which 22 bohrs was used. The correlation energy of the ACFDT-based methods was evaluated with a total of 8 imaginary frequency points and 20 eigenvalues per valence electron for H2 ($N_{\rm eig}$=40), 40 for LiH ($N_{\rm eig}$=80), and 10 for every remaining system ($N_{\rm eig}$=80 for FH, H2O, NH3, CH4; $N_{\rm eig}$=100 for N2). Results obtained for the structural properties are presented in Tab. 1, and the harmonic vibrational frequencies are presented in Tab. 2. The mean absolute percentage error (MAPE) indicated in these two tables is determined with respect to selected reference CCSD(T) results. They are illustrated in Figs. 4-5.
@@
First of all, we note that the results obtained with self-consistent RPA are very similar to the results from non-self-consistent RPA. The RPA[PBE] approach, therefore, represents an interesting alternative to fully self-consistent RPA given its cheaper computational cost. Although we do not have access to self-consistent RPAx results, we expect a very similar behavior for this method. For HF, we see slightly larger differences between EXX and EXX[PBE], most likely due to a larger dissimilarity between EXX and PBE orbitals. Despite that, the differences in terms of both structural properties and vibrational frequencies remain very small.
If we now compare the performance of the different approximations with respect to the most recent CCSD(T) results, we see that RPA generally overestimates the bond distances and underestimates the vibrational frequencies of the ”bond stretching” type. Nonetheless, the RPA method systematically improves the results when comparing it to PBE, lowering the MAPE error from 2.72% to 2.11%, on the average. A larger improvement is seen for the vibrational modes of ”bending” type, present in H2O, NH3 and CH4. The RPA reduces the 4.14% error in PBE to 0.75%, despite the fact that the bond angles are slightly worse in RPA as compared to PBE. The molecular geometries predicted by our implementation are in good agreement with the all-electron RPA calculations previously reported in the literature using Gaussian basis sets [30, 28, 29].
@@ V.2 Diamond, silicon, and germanium
We will now study the performance of the different ACFDT methods on solids. Let us start with the carbon diamond phase. Despite the simplicity of this system, which only has one degree of freedom - its lattice parameter - an accurate theoretical determination of the optical phonon at $\Gamma$ is still missing. Calculations performed at the DMC level have predicted a harmonic frequency at 1374-1384 cm-1 [71, 72], which overestimates the experimental frequency at 1334 cm-1 [74]. If the accuracy of the DMC calculations is within the stochastic error bar (4-8 cm-1), these results would suggest a quite large anharmonic renormalization, which disagrees with previous estimations done by total energy finite-difference calculations up to fourth order [79] and from path integral molecular dynamics calculations [72], which estimate a correction of -17.3 cm-1 and -13.7 cm-1, respectively.
Given this inconsistency, we have performed highly converged calculations with RPA, RPAx, RPAx(1), and AC-SOSEX. We have also applied the PBE0 hybrid functional with a fraction of exact-exchange optimized through RPA (17%) and RPAx (34%) [18, 38]. The lattice parameter, calculated for some of the methods, is presented in Tab. 3. The harmonic optical $\Gamma$-phonon, calculated at the experimental lattice constant ($a$ = 3.567 Å [74]), is presented in Tab. 4.
We first note that the RPA lattice parameter is in seemingly perfect agreement with the experimental value. However, subtracting the zero-point vibrational energy (ZPE) contribution from the experimental extrapolated zero temperature result gives $a$ = 3.553 Å, as shown in Ref. [73]. By comparing this value to our theoretical estimates, we find instead RPAx[PBE] to be in excellent agreement with experiments. Looking at the other methods, we see that PBE is close to RPA, while the RPA-optimized hybrid PBE0(17%) agrees well with RPAx[PBE]. On the other hand, the RPAx-optimized hybrid returns a too small estimate for $a$.
@@
The impact of self-consistency turns out to be negligible in RPA. Self-consistent RPA gives a frequency 1 cm-1 lower than RPA[PBE] in the $\times3$$\times$3 and $\times4$$\times$4 $\mathbf{k}$-point-grid calculations. We, therefore, did not pursue self-consistent RPA calculations with larger grids. Similarly, different starting points for RPAx (PBE, EXX) also showed very small differences in the frequency.
The RPA predicts a harmonic frequency at 1342 cm-1, which is in good agreement with Ref. [40], and close to the experimental value at 1334 cm-1. Adding the anharmonic correction of 17 cm-1, found in earlier works, gives us an experimental harmonic reference at 1351 cm-1. The RPA thus underestimates experiments by 9 cm-1, which is less than 1%. This can be compared to the PBE error of 4.5%, or 61 cm-1. The harmonic RPAx[PBE] value is at 1362 cm-1, which corresponds to an increase of 20 cm-1 with respect to RPA, and of an 11 cm-1 overestimation with respect to the harmonic experimental frequency. Clearly, the RPA and RPAx calculations produce a phonon frequency in closer agreement with experiments than DMC, which overestimates by 20-40 cm-1, taking into account error bars. If we consider the optimized hybrid functionals, we see that PBE0(17%) is only slightly below RPA, and PBE0(34%) only slightly above RPAx. This agreement tends to indicate that these methods work quite well in approximating the performance of the method that was used to optimize them, at least for the phonon frequencies.
@@
In order to understand the impact of the choice of functional for determining the anharmonic shift, we have followed the same finite-difference approach as in Ref. [75] but with the RPAx-optimized hybrid functional, which we found to give similar results to RPAx for the phonon frequency. Using this method, we find the correction to be about -14 cm-1, only slightly smaller than that of PBE. As such, the inclusion of exact exchange, crucial for the phonon frequency, appears to decrease the size of the anharmonic correction, and gives us a new experimental harmonic reference of 1348 cm-1, which slightly decreases the difference with respect to RPA and slightly increases the difference with respect to RPAx.
@@
Finally, the alternative RPAx resummations, RPAx(1) and AC-SOSEX, give us a spread of 20 cm-1 of possible values when including the exact-exchange kernel. RPAx(1) predicts a harmonic frequency at 1348 cm-1, which perfectly agrees with the experimental value, while AC-SOSEX predicts a value at 1368 cm-1, which is 6 cm-1 above the RPAx result. On the molecular systems studied in Sec. V.1, RPAx(1) systematically underestimated the vibrational frequencies, while AC-SOSEX overestimated them. RPAx also slightly overestimated the frequencies. If we assume that this trend holds for solids, the true frequency should lie somewhere between RPAx(1) and RPAx, i.e., in the interval 1348-1362 cm-1. This implies an overestimation with respect to experiments of just a few cm-1. At the same time, it is reasonable to expect technical errors, such as the pseudopotential approximation, to be of the order 5 cm-1.
@@
Thanks to our force implementation in RPA, we have also calculated the full phonon dispersion. Since there is no impact of self-consistency, we have only considered RPA[PBE] to save computational cost. We have used the same computational setup as before, only decreasing the number of eigenvalues per valence electron from 10 to 5, as the change in $\Gamma$-point frequency was found minimal, only affecting the frequency by 2 cm-1 on the $\times6$$\times$6 grid. We then constructed a [222] diamond supercell containing 16 atoms. For the choice of sampling, we used the same setting as our most advanced $\Gamma$-point calculation, which corresponds to a $\times3$$\times$3 $\mathbf{k}$-point grid. The RPA phonon dispersion calculated on the experimental unit cell is given in Fig. 6, together with the results in PBE, PBE0(17%), and PBE0(34%). As can be seen, it is mostly the optical branches that are affected by the choice of functional.
@@
The second material that we have considered in this study is the diamond phase of silicon. On this system, we have performed a similar study as for C. Previous works have shown that the lattice constant is rather sensitive to the pseudopotential approximation [38]. We have, therefore, first verified that the phonon frequencies do not show a similar sensitivity when evaluated at the experimental lattice constant. Using different pseudopotential libraries [82, 83] we found discrepancies of maximum 2 cm-1 for RPA and RPAx, and 3 cm-1 for PBE. For this study, we have, therefore, continued to use the ONCV library.
@@
We then computed the optical $\Gamma$-phonon using PBE, the RPA-optimized hybrid PBE0(29%) [38], RPA, RPAx, RPAx(1), and AC-SOSEX. For the PBE, PBE0(29%), and RPAx calculations, we used a kinetic energy cutoff of 80 Ry. For RPA 60 Ry was sufficient. For the correlation energy, we used a grid of 8 imaginary frequency points and determined 60 eigenvalues of the response function in both RPA and RPAx. The total energy finite-difference calculations were obtained with mixed $\mathbf{k}$-point grids. The EXX energy was converged on a $\times8$$\times$8 $\Gamma$-centered grid, while the correlation energy was converged on a shifted $\times4$$\times$4 grid. The value of the $\Gamma$ phonon frequency for all methods is listed in Tab. 5. The RPA[PBE] predicts a frequency at 522 cm-1 in very good agreement with the experimental value at 523 cm-1 [78]. Effects of self-consistency in RPA was again negligible. The RPAx frequency lies at 531 cm-1, RPAx(1) at 527 cm-1 and AC-SOSEX at 536 cm-1. The anharmonic shift has been estimated to -3.5 cm-1 [79, 75]. From our analysis on diamond we expect this value to slightly decrease using an improved functional. Comparing to the estimated harmonic experimental frequency at 526 cm-1, the RPA underestimates by 3 cm-1, RPAx/AC-SOSEX overestimate by 5/10 cm-1, and RPAx(1) is exactly on top of the experimental value. This trend is the same as seen in diamond. For PBE and PBE0(29%), we used a $\times8$$\times$8 $\Gamma$-centered grid. PBE underestimates the frequency by 12 cm-1 and PBE0(29%) overestimates it by 10 cm-1.
@@
The phonon dispersion is presented in Fig. 7. Similarly to diamond, it has been acquired using the forces on a [222] supercell. A $\times4$$\times$4 grid was considered for both PBE and PBE0(29%). For RPA[PBE] we used a $\times3$$\times$3 sampling. At this sampling, the $\Gamma$-mode agrees, within 2 cm-1, with our fully converged total-energy finite difference calculation above. Going from $\times2$$\times$2 to $\times3$$\times$3 resulted in very small changes, mainly around the $L$-point and for the first acoustic branch around the $X$-point. We are, therefore, confident that the phonon dispersion is converged in RPA[PBE] with a $\times3$$\times$3 grid on the supercell.
@@
The $\Gamma$-phonon was then evaluated at experimental lattice parameter, with the total-energy finite-difference approach, for RPA[OEP(33%)] and self-consistent RPA. At a $\Gamma$-centered $\times4$$\times$4 $\mathbf{k}$-point-grid this gave a difference of 2 cm-1 in the phonon frequency. We, therefore, continued using OEP(33%) as a starting point for RPAx, RPAx(1) and AC-SOSEX. We found a quick convergence with respect to the number of eigenvalues of the response function. Results were well-converged already with a total of 20-40 eigenvalues. On the other hand, due to the smaller gap in Ge, we needed to include up to 12 imaginary frequency points. A plane-wave cut-off of 60 Ry was sufficient. For the $\mathbf{k}$-point convergence we used a similar setup as in silicon: $\times8$$\times$8 for the EXX part of the energy and a $\times4$$\times$4 shifted grid for the correlation energy. The results for the $\Gamma$-phonon are presented in Tab. 6. We have also included the frequency in PBE0(33%) and PBE, where we used a $\times8$$\times$8 $\mathbf{k}$-point grid and a 80 Ry plane-wave cutoff.
The anharmonic correction was estimated also for Ge [75] giving a value of -1.2 cm-1. Taking this into account, the experimental harmonic phonon is at 305 cm-1. The RPA[OEP(33%)] value is exactly on top of this value. We see that RPA corrects the PBE frequency by 10 cm-1. The effect of exact-exchange kernel is rather small, increasing the RPA frequency by 4 cm-1 in RPAx, 1 cm-1 in RPAx(1), and 5 cm-1 in AC-SOSEX. Thus, all ACFDT-based methods predict a value in good agreement with experiments. The PBE0(33%) functional gives phonon at 310 cm-1, which, similar to C and Si, slightly overestimates the RPA result.
We have not included the Ge dispersion plot in RPA due to the slow $\mathbf{k}$-point convergence of the phonons around the $L$-point, which would necessitate larger grids than $\times3$$\times$3 on the [222] supercell. In addition, a self-consistent calculation would be needed which multiplies the cost by the number of iterations required to converge the KS orbitals.

### v2

@@ IV Implementation and testing
Given the good quality found for our forces on this first test, we have then proceeded to determine their quantitative accuracy by comparing our analytical force calculations to the total energy finite-difference estimations. For this analysis, two different systems have been considered: the H2O molecule and the wurtzite phase of the boron nitride solid ($\omega$-BN, space group $P6_{3}/mmc$). For the H2O molecule, calculations were performed using a 120 Ry cutoff for the kinetic energy, and 20 bohrs for the simulation cell size. With these settings, the PBE structure and PBE vibrational frequencies were well converged when using the forces. For the RPA calculations, we used the same plane-wave cutoff and cell size. For the evaluation of the correlation components, we used 10 eigenvalues per valence electron ($N_{\rm eig}$=80), and 12 imaginary frequency points. Starting from the RPA equilibrium geometry (see Tab. I), we then varied one of the O-H bond length around 0.96 Å, in steps of 0.01 Å. The total energy finite-difference estimations were performed using the highly accurate five-point stencil derivation formula. For the $\omega$-BN solid, we first optimized the unit cell using the PBE functional and then displaced one B3+ cation along the $\mathbf{c}$ crystal axis. For these test calculations, we used a cutoff of 100 Ry, and a $\times2$$\times$2 k-point grid. For RPA, we used a grid made of 10 imaginary frequency points, determined 8 eigenvalues per valence electron ($N_{\rm eig}$=128), and used a $\times2$$\times$2 q-point grid. The error analyses for H2O and $\omega$-BN are presented in panels a) and b) of Fig. 3, respectively.
@@
In panel a), we note that the largest error on the force is found for PBE, with a value of $2\!\times\!10^{-4}$ Ry/bohr, substantially above any of the other approximations considered. This error reduces, however, by increasing the energy cutoff, either on the plane waves or the density. At the level of HF, we instead note an almost perfect agreement between the analytical evaluation of the force and the numerical estimate from the total energy, with a difference between the two predictions below $10^{-6}$ Ry/bohr. The EXX method displays, at most, an error of $1.6\!\times\!10^{-5}$ Ry/bohr across the whole range of distances monitored. As discussed in Ref. [42], the oscillations and the overall error can be dampened further by increasing the kinetic energy cutoff and by tightening the self-consistent convergence conditions. If we compare the effect of self-consistency in the calculation, we see that the error found at the HF[PBE] level is larger than for HF and EXX. Despite this, the overall agreement remains of good quality, with a maximum error below $6.5\!\times\!10^{-5}$ Ry/bohr. A similar evolution is observed for the RPA method. At self-consistency, the error on RPA forces is found below $1\!\times\!10^{-5}$ Ry/bohr, similar to EXX. Without self-consistency, the evolution of RPA[PBE] matches HF[PBE], with a maximum error of $5.3\!\times\!10^{-5}$ Ry/bohr. Overall, the errors found are sufficiently small for exploiting the forces in a geometry relaxation procedure, given the 0.001 Ry/a0 default threshold set on each individual force within QE.
@@
Finally, in both panels a) and b), we note that if we omit the extra xc force terms (Eq. (18)) for EXX and RPA, the methods denoted EXX $\{-{\bf F}^{\rm scf\text{-}EXX}_{\rm x}\}$ and RPA $\{-{\bf F}^{\rm scf\text{-}RPA}_{\rm x}\}$, respectively, the error in the analytical evaluation of the force is as large as 0.1-0.01 Ry/a0. This further confirms the importance of including these terms for any approximation that relies on the OEP method.
@@ V Results and discussion
Below we present our results for a set of molecules (H2, LiH, HF, N2, H2O, NH3, and CH4), and periodic solids (C, Si, and Ge).
@@ V.1 Molecules
We have evaluated the performance of the ACFDT-based methods with respect to HF, PBE, and reference CCSD(T) calculations on a set of covalently bonded molecules: H2, LiH, HF, N2, H2O, NH3, and CH4. For every system considered, a kinetic energy cutoff of 100 Ry has been used, and a cubic simulation cell of dimensions 20 bohrs has been used except for LiH, N2, and CH4, for which 22 bohrs was used instead. The correlation energy of the ACFDT-based methods was evaluated with a total of 8 imaginary frequency points, and 20 eigenvalues per valence electron for H2 ($N_{\rm eig}$=40), 40 for LiH ($N_{\rm eig}$=80), and 10 for every remaining system ($N_{\rm eig}$=80 for HF, H2O, NH3, CH4, and $N_{\rm eig}$=100 for N2). Results obtained for the structural properties are presented in Tab. 1, and the harmonic vibrational frequencies are presented in Tab. 2. The mean absolute percentage error (MAPE) indicated in these two tables is determined with respect to selected reference CCSD(T) results. They are illustrated in Figs. 4 and 5.
@@
First of all, we note that the results obtained with self-consistent RPA are very similar to the results from non-self-consistent RPA. The RPA[PBE] approach, therefore, represents an interesting alternative to fully self-consistent RPA given its cheaper computational cost. Although we do not have access to self-consistent RPAx results, we expect a very similar behavior for this method. For HF, we see slightly larger differences between EXX and EXX[PBE], most likely due to a larger dissimilarity between the EXX and PBE orbitals. Despite that, the differences in terms of both structural properties and vibrational frequencies remain very small.
If we now compare the performance of the different approximations with respect to the most recent CCSD(T) results, we see that RPA generally overestimates the bond distances and underestimates the vibrational frequencies of the ”bond stretching” type. Nonetheless, the RPA method systematically improves the results when comparing it to PBE, lowering the MAPE from 2.72% to 2.11%, on the average. A larger improvement is seen for the vibrational modes of ”bending” type, present in H2O, NH3, and CH4. The RPA reduces the 4.14% error in PBE to 0.75%, despite the fact that the bond angles are slightly worse in RPA as compared to PBE. The molecular geometries predicted by our implementation are in good agreement with the all-electron RPA calculations previously reported in the literature using Gaussian basis sets [31, 29, 30].
@@ V.2 Diamond, silicon, and germanium
We will now study the performance of the different ACFDT methods on solids. Let us start with the carbon diamond phase. Despite the simplicity of this system, which only has one degree of freedom - its lattice parameter - an accurate theoretical determination of the optical phonon at $\Gamma$ is still missing. Calculations performed at the DMC level have predicted a harmonic frequency at 1374-1384 cm-1 [74, 75], which overestimates the experimental frequency at 1334 cm-1 [77]. If the accuracy of the DMC calculations is within the stochastic error bar (4-8 cm-1), these results would suggest a quite large anharmonic renormalization, which disagrees with previous estimations done by total energy finite-difference calculations up to fourth order [82] and by PBE path integral molecular dynamics calculations [75], which estimate a correction of -17.3 cm-1 and -13.7 cm-1, respectively.
Given this inconsistency, we have performed highly converged calculations with RPA, RPAx, RPAx(1), and AC-SOSEX at the harmonic level. In order to estimate the anharmonic shift as consistently as possible, yet in a computationally feasible way, we have optimized the fraction of exact exchange in the PBE0 hybrid functional, according to the procedure in Refs. [18, 39]. In RPA we find a fraction of 17% and in RPAx 34%. The lattice parameter, calculated for some of the methods, is presented in Tab. 3. The harmonic optical $\Gamma$-phonon, calculated at the experimental lattice constant ($a$ = 3.567 Å [77]), is presented in Tab. 4.
We first note that the RPA lattice parameter is in seemingly perfect agreement with the experimental value. However, subtracting the zero-point vibrational energy (ZPE) contribution from the experimental extrapolated zero-temperature result gives $a$ = 3.553 Å, as shown in Ref. [76]. By comparing this value to our theoretical estimates, we find instead RPAx[PBE] to be in excellent agreement with experiments. Looking at the other methods, we see that PBE is close to RPA, while the optimized hybrid functionals underestimate the lattice constant with respect to the prediction of the underlying RPA/RPAx method.
@@
The impact of self-consistency turns out to be negligible in RPA. Self-consistent RPA gives a frequency 1 cm-1 lower than RPA[PBE] in the $\times3$$\times$3 and $\times4$$\times$4 $\mathbf{k}$-point-grid calculations. We, therefore, did not pursue self-consistent RPA calculations with larger grids. Similarly, different starting points for RPAx (such as PBE and EXX) also showed very small differences in the frequency.
The RPA predicts a harmonic frequency at 1342 cm-1, which is in good agreement with Ref. [41], and close to the experimental value at 1334 cm-1. Adding the PBE anharmonic correction of -17 cm-1, found in earlier works, gives us an experimental harmonic reference at 1351 cm-1. The RPA thus underestimates experiments by 9 cm-1, which is less than 1%. This can be compared to the PBE error of 4.5%, or 61 cm-1. The harmonic RPAx[PBE] value is at 1362 cm-1, which corresponds to an increase of 20 cm-1 with respect to RPA, and to an 11 cm-1 overestimation with respect to the harmonic experimental frequency. Clearly, the RPA and RPAx calculations produce a phonon frequency in closer agreement with experiments than DMC, which overestimates by 20-40 cm-1, taking into account error bars. If we consider the optimized hybrid functionals, we see that PBE0(17%) is only slightly below RPA, and PBE0(34%) only slightly above RPAx. This agreement suggests that these methods work quite well in approximating the performance of the method that was used to optimize them.
@@
In order to understand the impact of the choice of functional for determining the anharmonic shift, we have followed the same finite-difference approach as in Ref. [78] but with the RPAx-optimized hybrid functional, which we found to give similar results to RPAx for the phonon frequency. Using this method, we find the correction to be -14.2 cm-1, only slightly smaller than that of PBE, -17.4 cm-1. As such, the inclusion of exact exchange, crucial for the phonon frequency, appears to decrease the size of the anharmonic correction, and gives us a new experimental harmonic reference of 1348 cm-1, which slightly decreases the difference with respect to RPA and slightly increases the difference with respect to RPAx.
@@
Finally, the alternative RPAx resummations, RPAx(1) and AC-SOSEX, give us a spread of 20 cm-1 of possible values when including the exact-exchange kernel. RPAx(1) predicts a harmonic frequency at 1348 cm-1, which perfectly agrees with the experimental value, while AC-SOSEX predicts a value at 1368 cm-1, which is 6 cm-1 above the RPAx result. On the molecular systems studied in Sec. V.1, RPAx(1) systematically underestimated the vibrational frequencies, while AC-SOSEX overestimated them. RPAx also slightly overestimated the frequencies. If we assume that this trend holds for solids, the true frequency should lie somewhere between RPAx(1) and RPAx, i.e., in the interval 1348-1362 cm-1. This implies an overestimation with respect to experiments of just a few cm-1. At the same time, it is reasonable to expect technical errors, such as the pseudopotential approximation, to be on the order of 5 cm-1.
@@
Thanks to our force implementation in RPA, we have also calculated the full phonon dispersion. Since there is no impact of self-consistency, we have only considered RPA[PBE] to save computational cost. We have used the same computational setup as before, only decreasing the number of eigenvalues per valence electron from 10 to 5, as the change in $\Gamma$-point frequency was found to be minimal, only affecting the frequency by 2 cm-1 on the $\times6$$\times$6 grid. We then constructed a [222] diamond supercell containing 16 atoms. A supercell allows for a precise determination of the $\mathbf{q}$-points that are commensurate with its size, in our case, $\Gamma$, X, and L. The points between these high-symmetry $\mathbf{q}$-points are, however, approximated through a Fourier interpolation. For the choice of sampling, we used the same setting as our most advanced $\Gamma$-point calculation, which corresponds to a $\times3$$\times$3 $\mathbf{k}$-point grid on the supercell. The RPA phonon dispersion calculated on the experimental unit cell is given in Fig. 6, together with the results in PBE, PBE0(17%), and PBE0(34%). As can be seen, it is mostly the optical branches that are affected by the choice of functional.
@@
The second material that we have considered in this study is the diamond phase of silicon. On this system, we have performed a similar study as for C. Previous works have shown that the lattice constant is rather sensitive to the pseudopotential approximation [39]. We have, therefore, first verified that the phonon frequencies do not show a similar sensitivity when evaluated at the experimental lattice constant. Using different pseudopotential libraries [85, 86], we found discrepancies of maximum 2 cm-1 for RPA and RPAx, and 3 cm-1 for PBE. For this study, we have, therefore, continued to use the ONCV library.
@@
We then computed the optical $\Gamma$-phonon using PBE, the RPA-optimized hybrid PBE0(29%) [39], RPA, RPAx, RPAx(1), and AC-SOSEX. For the PBE, PBE0(29%), and RPAx calculations, we used a kinetic energy cutoff of 80 Ry. For RPA, 60 Ry was sufficient. For the correlation energy, we used a grid of 8 imaginary frequency points and determined 60 eigenvalues of the response function in both RPA and RPAx. The total energy finite-difference calculations were obtained with mixed $\mathbf{k}$-point grids. The EXX energy was converged on a $\times8$$\times$8 $\Gamma$-centered grid, while the correlation energy was converged on a shifted $\times4$$\times$4 grid. The value of the $\Gamma$ phonon frequency for all methods is listed in Tab. 5. The RPA[PBE] predicts a frequency at 522 cm-1, in very good agreement with the experimental value at 523 cm-1 [81]. Effects of self-consistency in RPA was again negligible. The RPAx frequency lies at 531 cm-1, RPAx(1) at 527 cm-1, and AC-SOSEX at 536 cm-1. Using the same approach as before on C, we estimate the anharmonic shift of Si to be -3.1 cm-1 with PBE and -2.9 cm-1 with the optimized PBE0(29%) hybrid, in good agreement with the value of -3.5 cm-1 already presented in the literature [78]. Compared to the estimated harmonic experimental frequency at 526 cm-1, the PBE underestimates the frequency by 12 cm-1, the RPA underestimates by 3 cm-1, RPAx/AC-SOSEX overestimate by 5/10 cm-1, and RPAx(1) is exactly on top of the experimental value. This trend is the same as seen in diamond.
@@
The phonon dispersion is presented in Fig. 7. Similarly to diamond, it has been acquired using the forces on a [222] supercell. A $\times4$$\times$4 $\mathbf{k}$-point grid was considered for both PBE and PBE0(29%). For RPA[PBE], we used a $\times3$$\times$3 sampling. At this sampling, the $\Gamma$-mode agrees, within 2 cm-1, with our fully converged total energy finite difference calculation presented above. Going from $\times2$$\times$2 to $\times3$$\times$3 resulted in very small changes in the dispersion plot, mainly around the L-point and for the first acoustic branch around the X-point. We are, therefore, confident that the phonon dispersion, in the vicinity of the high-symmetry $\mathbf{q}$-points, is converged in RPA[PBE] with a $\times3$$\times$3 grid on the supercell.
@@
The $\Gamma$-phonon was then evaluated at the experimental lattice parameter, with the total energy finite-difference approach, for RPA[OEP(33%)] and self-consistent RPA. At a $\Gamma$-centered $\times4$$\times$4 $\mathbf{k}$-point-grid, we found a difference of only 2 cm-1 in the phonon frequency between the two methods. We, therefore, continued using OEP(33%) as a starting point for RPAx, RPAx(1), and AC-SOSEX. We found a quick convergence with respect to the number of eigenvalues of the response function. Results were well-converged already with a total of 20-40 eigenvalues. On the other hand, due to the smaller gap in Ge, we needed to include up to 12 imaginary frequency points. A plane-wave cutoff of 60 Ry was sufficient. For the $\mathbf{k}$-point convergence, we used a similar setup as in silicon: $\times8$$\times$8 for the EXX part of the energy and a $\times4$$\times$4 shifted grid for the correlation energy. The results for the $\Gamma$-phonon are presented in Tab. 6. We have also included the frequency in PBE0(33%) and PBE, where we used a $\times8$$\times$8 $\mathbf{k}$-point grid and a 80 Ry plane-wave cutoff.
The anharmonic correction was estimated also for Ge in Ref. [78], giving a value of -1.2 cm-1. Taking this into account, the experimental harmonic phonon is at 305 cm-1. The RPA[OEP(33%)] value is exactly on top of this value. We see that RPA corrects the PBE frequency by 10 cm-1. The effect of the exact-exchange kernel is rather small, increasing the RPA frequency by 4 cm-1 in RPAx, 1 cm-1 in RPAx(1), and 5 cm-1 in AC-SOSEX. Thus, every ACFDT-based method predicts a value in good agreement with experiments. The PBE0(33%) functional finds the optical phonon at 310 cm-1, which, similar to C and Si, slightly overestimates the RPA result.
We have not included the Ge dispersion plot in RPA due to the slow $\mathbf{k}$-point convergence of the phonons around the L-point, which would necessitate larger grids than $\times3$$\times$3 on the [222] supercell. In addition, a self-consistent calculation would be needed, which multiplies the cost by the number of iterations required to converge the KS orbitals.